A galvanic cell is set up from a zinc bar weighing 100g and 1.0 litre of 1M CuSO4 solution. How long would the cell run if it is assumed to deliver a deliver a stedy current of 1 amp:
(atomic mass of Zn=65)
a]1.1 hr.             b]53.6 hr.

Grade:

1 Answers

Arun
4 years ago
Moles of zinc reduced
= 100/65
= 1.53 mol
Here, CuSO4 is limiting reagent. 
Cu2+ + 2e- ---> Cu
That is to completely dissolve 1 mol of Cu, 2 moles of electrons are required.
Time for which cell runs is -
t = Q / I
t = 2 × 96500 / 1
t = 193000 s
t = 53.61 hrs
Therefore, cell will run for 53.61 hrs.